Attendees: Bruce Glagola, Bruno Fieramosca, Clay White, Nena Moonier, Wendy VanWingeren, Dean Wyncott, and Patti Pedergnana
- Tomorrow is the first planning/scheduling meeting for shutdown in building 401 room B4100.Wendy will attend and Bruno will take the 0101 pager at during that time.
- There have been meetings regarding Tecknit leaving the APS.The users at Sectors 8, 16, and 35 are working with John Wright to get their work done before April 30t. One new idea for Tecknit is to do batch work SSRO’s.
- Today is Tech Tuesday, so Bruno will visit 431, Clay 433, and Dean 438.
- The new RSS policy and procedure is not complete yet.Julie Cross will visit with our group when it is done.
- ESAF will be updated to include the start time and end time on the EA form.The host name will also be added.
- There is a routine radiation survey tomorrow.
- Dean is catching up on the electrical legends in the panels.
- Sector 15 has explosive samples to be packaged and shipped out.
- There are three radioactive sample experiments to be set up tomorrow, one at Sector 13 and two at Sector 16.
- Shane and Patti are working on the UES homepage getting moved to Drupal.
- Bruce approved the Techniplast request for the biosafety cabinet certifications.
- There is a BSL-2 experiment starting at Sector 23 on April 2nd.
- The CCWP is in progress for the 8-BM removal of the outboard wall.The project start date was moved up to April 20th.
- If there is a vacuum issue at 4-ID-C, it may actually be a glitch in the system.Beamline personnel (Freeland, Keavney) can reset this remotely.FCs will see no vacuum fault on EPICS even though the beamline will not be able to run..